Va:

A local hamburger shop sold a combined total of 478 hamburgers and cheeseburgers on Sunday. There were 72 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Sunday?

jhonyy9:

let the hamburgers be x and cheeseburgers y x +y = 478 x = y -72 solve this system for x and y

jhonyy9:

@Laylalyssa opinion pls. ? ty

jhonyy9:

@Convert opinion pls. ty

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 let the hamburgers be x and cheeseburgers y x +y = 478 x = y -72 solve this system for x and y \(\color{#0cbb34}{\text{End of Quote}}\) Yes, I believe that'd be correct. Good job jhonyy! (: